# The non-reciprocal Ising model

### v2

@@ Abstract
Systems with non-reciprocal interactions generically display time-dependent states. These are routinely observed in finite systems, from neuroscience to active matter, in which globally ordered oscillations exist. However, the stability of these uniform non-reciprocal phases in noisy spatially-extended systems, their fate in the thermodynamic limit, and the criticality of the corresponding phase transitions are not fully understood. Here, we address these questions by introducing a non-reciprocal generalization of the Ising model and study its critical behavior by means of numerical and analytical approaches. While the mean-field equations predict three stable homogeneous phases (disordered, ordered and a time-dependent swap phase), our large scale numerical simulations reveal a more complex picture. Static order is destroyed in any finite dimension due to the growth of rare droplets unless the symmetry between the two spins types is broken triggering a stabilizing droplet-capture mechanism. The swap phase is destroyed by fluctuations in two dimensions through the proliferation of spiral defects, but stabilized in three dimensions where non-reciprocity changes the critical exponents from Ising to XY, thus giving rise to a robust spatially-distributed clock.
@@
Non-reciprocal interactions naturally arise in out-of-equilibrium systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] ranging from neuroscience [23, 24, 25] and social networks [26, 27] to ecology [28, 29, 30] and open quantum systems [31, 32, 33]. A generic feature of these systems is the emergence of many-body limit cycles – time-dependent states arising from non-mutual interactions between constituents. Such states are routinely observed in experiments and simulations of finite systems. A single limit cycle oscillator subject to noise eventually forgets the initial phase [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]: its temporal correlations decay with time very much like density fluctuations decay with spatial separation in a liquid (Fig. 1a, left). In many-body systems, however, coherent oscillations may be restored over arbitrarily long periods [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 50, 58, 59, 49, 60, 61, 62, 63, 64]: temporal correlations persist (without any periodic drive externally imposing a phase) very much like spatial correlations do in a crystal (Fig. 1a, right). In this Letter, we ask: Can many-body limit cycles survive fluctuations in spatially-extended, locally-coupled non-reciprocal systems of arbitrary size? If so, what are the critical exponents of the resulting phase transitions? We address these questions by introducing a non-reciprocal generalization of what is perhaps the most paradigmatic statistical mechanical system: the Ising model.
@@
Dynamics without global optimisation.— In equilibrium, the dynamics of Ising spins is described by kinetic Ising models in which spins tend to minimize a global potential, and hence interact reciprocally [65, 66]. To account for non-reciprocal interactions, we instead assume that each spin tends to minimize its own selfish energy , and choose the probability of flipping at each time step to be given by the Glauber rule
@@
where is the change in selfish energy between both configurations, is the temperature, and the Boltzmann constant. Since we define different potential functions for each spin, the dynamics cannot be derived from a single potential function and detailed balance is broken. This formulation encompasses several non-reciprocal dynamics used among the physical sciences [67, 68, 69, 70, 71, 72, 73, 74, 75, 76].
@@
The non-reciprocal Ising model.— Systems with non-reciprocal couplings do not always exhibit macroscopic oscillations. These tend to occur, when a small number of species have competing goals. Correspondingly, we split the spins into two species, labeled by Greek indices , located on sites of a -dimensional cubic lattice of linear size (Fig. 1b). Spins of the same species tend to align with their neighbors, while spins of different species interact in a non-reciprocal manner: spins tend to align with spins , whereas spins tend to anti-align with spins . This is captured by the selfish energy
@@
with , where the first term captures the intra-species interactions with the sum running over nearest neighbors of , and the second term captures the inter-species interactions where summation over species indices is implied ( is the Levi-Civita symbol). We have assumed for simplicity that inter-species interactions only occur on-site and are purely non-reciprocal (reciprocal on-site interactions are considered in [77]).
@@
Mean-field equation.— Within the mean-field approximation and upon rescaling of time and space, the dynamics of the average magnetization is [77]
@@
The phase diagram in Fig. 1c shows the stable homogeneous solutions of Eq. (3) as a function of the couplings and . When non-reciprocal interactions are turned off (), a pitchfork bifurcation at (red point in Fig. 1c) separates a disordered phase (in blue) from a ferromagnetic phase (in green), as in the equilibrium Ising model. When non-reciprocal interactions are present (), a time-dependent oscillatory state that we dub “swap phase” arises (in pink). This limit cycle state where both species flip their magnetizations repeatedly is separated from the disordered state by a Hopf bifurcation at (green line), and from the ferromagnetic phase by another bifurcation known as a saddle-node on an invariant circle (SNIC) bifurcation [78, 79] (blue line). The shape of the limit cycle in () space evolves from a circle near the Hopf bifurcation to a square near the SNIC bifurcation [77]. The oscillation period remains finite at the Hopf bifurcation, and it diverges at the SNIC bifurcation [78].
Monte-Carlo simulations.— To go beyond the mean-field description, we perform large-scale Monte-Carlo (MC) simulations of the non-reciprocal Ising model in both two and three dimensions (see [77] for details on the update rules; we did not observe qualitative differences with other update rules). Unless mentioned otherwise, we initialize the system in an ordered state where all spins of the same species are either up or down, and let the system evolve until a steady-state is reached.
The qualitative results of our simulations are summarized in Fig. 1d-e. In 2D, any amount of non-reciprocity destroys all order. In 3D, our simulations suggest that the swap phase survives the thermodynamic limit, while the static ferromagnetic phase is eventually destroyed by any amount of non-reciprocity.
@@
and the phase space angular momentum 111The sign of the phase space angular momentum in Eq. (5) is chosen such that it is positive when is positive. is related to the entropy production [121] close to the Hopf bifurcation.
where is the angle on the plane (Fig. 2a), are the total magnetizations and the average is over time and realizations. The synchronization order parameter is zero if the system is disordered and nonzero in both the swap and static-order phases, while is zero in the disordered and static-order phases and non-zero in the swap phase.
Destruction of the swap phase in 2D by spiral defects.— In small 2D systems, simulations show states in which and oscillate in time (like in the mean-field swap phase) with a large amplitude. However, oscillations are irregular, and typically occur through nucleation of droplets of opposite magnetization by the “unsatisfied” species, in alternating order (see Movie 1). As the size increases, it becomes apparent that this oscillatory state is a transient, eventually destabilized by proliferation of spiral defects similar to the rotational-symmetric spirals observed in the complex Ginzburg-Landau equation [81, 82, 83, 47, 84, 85, 86, 87, 88, 89, 90], but whose shape has 4-fold rotational symmetry (four “arms”). This is evidenced in Fig. 2b-c, which shows the emergence of defects in the angle and the resulting drop in the magnetizations and after an oscillatory transient, along with snapshots of at different stages (see also Movie 1).
@@
Existence of a stable swap phase in 3D and critical exponents.— In 3D, we find that a stable swap phase does exist. Its behavior is demonstrated in Fig. 3a where we show a kymograph of the -field of a single 1D row in the 3D lattice as well as plot the total magnetizations as a function of time (see also Movie 2). The oscillations are spatially homogeneous and and have a fixed period and phase shift. The coherence time of and diverges with system size as (Fig. 3b), in agreement with a temporal crystal behavior (Fig. 1a) [91]. In the thermodynamic limit, coherent oscillations persist indefinitely despite the presence of noise.
@@
Figure 3c shows that the phase-space trajectories become circular when approaching the phase transition line, supporting a Hopf behavior ([92]). Renormalization group studies based on the -expansion suggest that the field theory associated with the Hopf bifurcation in a spatially extended system has the same critical exponents as the XY model [93, 94]. Using finite-size scaling [77] and setting we determine the critical exponents of the correlation length, susceptibility, and order parameter: , , and , which are in good agreement with 3D XY critical exponents [95, 96], more so than with 3D Ising critical exponents [95] corresponding to the case, see Fig. 3f.
For a fixed , the critical separating disorder from swap depends on , in contrast with the mean-field prediction, in which is independent of (Fig. 1c). Above some critical value, the phase transition loses its second-order-like nature (see region in Fig. 3d where becomes discontinuous), apparently due to scroll waves (3D analog of spiral waves) [97, 98, 99, 100, 101] [77] that destabilize the swap phase. Within the size limitations of our simulations, it is unclear what part of the swap phase is destroyed by scroll waves in the thermodynamic limit and what type of phase transition is associated with their appearance (Fig. 1e).
@@
Stability of static-order phase and droplet growth in any finite dimension.— In finite systems, a ferromagnetic-like state with static order is observed both in 2D and 3D. As system size increases, however, this static state is destabilized by nucleation of droplets (see Fig. 4a and Movie 4) that grow and flip the magnetization, in alternating order of A- and B-spins. As a result, the static-order phase is replaced by droplet-induced oscillations (different in nature from the homogeneous noisy oscillations close to the Hopf bifurcation) with oscillation period that converges to a finite value as system size increases (Fig 4b). This is further supported by Figs. 2e and 3e (right parts) showing that the -transition-point between static order and oscillations in finite systems (indicated by a sharp change in the slope of ) increases with size.
To see why non-reciprocity allows droplets to grow, in contrast with the equilibrium Ising model where they tend to shrink, assume first that the system is in the static-order phase, so most spins in both lattices are up [102]. Since the system is static, it can be mapped into two equilibrium Ising models (Fig. 4c) with opposite magnetic fields, with energy where for A-spins and for B-spins (see Eq. (2)). While sub-system A is in a stable state in the effective equilibrium system, sub-system B is in a metastable state, as it prefers a state with opposite magnetization. Sub-system then transitions to its equilibrium stable state by nucleating droplets larger than a critical size whose value depends on and . [103, 102]. After a droplet has expanded beyond , the stable sub-system becomes metastable and nucleates droplets, and so on (Fig. 4c). Crucially, for finite and , is finite, making the static-order phase unstable in any finite dimension. Note that when the symmetry between the spin species is broken, droplets do not expand at the same velocity, leading to droplet-capture and shrinkage that re-stabilizes the static-order phase (Fig. 4d and Movie 5). This can be achieved, for example, by modifying the inter-species interaction in the selfish energy, Eq. (2), to be where and (see Ref. [77] for details).
@@
We have shown that a non-reciprocal generalization of the 3D Ising model can act as a stable spatially-distributed clock characterized by well-defined critical properties reminiscent of time-crystals [110, 111, 112, 97, 113, 114, 115, 116]. While minimalistic, this model contains features arising in models of the human brain [67], opinion dynamics [117, 118], spinor BECs [119, 120, 31] and micromechanical oscillators [69], which can all be modeled by non-potential spin systems.
@@ Acknowledgements.
- [80] The sign of the phase space angular momentum in Eq. (5) is chosen such that it is positive when is positive. is related to the entropy production [121] close to the Hopf bifurcation.
@@
- [92] Yoshiki Kuramoto and Yoshiki Kuramoto. Chemical turbulence. Springer, 1984.

### v3

# Nonreciprocal Ising model

@@ Abstract
Systems with nonreciprocal interactions generically display time-dependent states. These are routinely observed in finite systems, from neuroscience to active matter, in which globally ordered oscillations exist. However, the stability of these uniform nonreciprocal phases in noisy spatially-extended systems, their fate in the thermodynamic limit, and the critical behavior of the corresponding phase transitions are not fully understood. Here, we address these questions by introducing a nonreciprocal generalization of the Ising model and study its phase transitions by means of numerical and analytical approaches. While the mean-field equations predict three stable homogeneous phases (disordered, ordered and a time-dependent swap phase), our large scale numerical simulations reveal a more complex picture. Static order is destroyed in any finite dimension due to the growth of rare droplets unless the symmetry between the two spin types is broken triggering a stabilizing droplet-capture mechanism. The swap phase is destroyed by fluctuations in two dimensions through the proliferation of spiral defects, but stabilized in three dimensions where nonreciprocity changes the critical exponents from Ising to XY, thus giving rise to a robust spatially-distributed clock.
@@
Nonreciprocal interactions naturally arise in out-of-equilibrium systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] ranging from neuroscience [23, 24, 25] and social networks [26, 27] to ecology [28, 29, 30] and open quantum systems [31, 32, 33]. A generic feature of these systems is the emergence of many-body limit cycles – time-dependent states arising from non-mutual interactions between constituents. Such states are routinely observed in experiments and simulations of finite systems. A single limit cycle oscillator subject to noise eventually forgets the initial phase [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]: its temporal correlations decay with time very much like density fluctuations decay with spatial separation in a liquid (Fig. 1a, left). In many-body systems, however, coherent oscillations may be restored over arbitrarily long periods [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 50, 58, 59, 49, 60, 61, 62, 63, 64]: temporal correlations persist (without any periodic drive externally imposing a phase) very much like spatial correlations do in a crystal (Fig. 1a, right). In this Letter, we ask: Can many-body limit cycles survive fluctuations in spatially-extended, locally-coupled nonreciprocal systems of arbitrary size? If so, what are the critical exponents of the resulting phase transitions? We address these questions by introducing a nonreciprocal generalization of what is perhaps the most paradigmatic statistical mechanical system: the Ising model.
@@
Dynamics without global optimization— In equilibrium, the dynamics of Ising spins is described by kinetic Ising models in which spins tend to minimize a global potential, and hence interact reciprocally [65, 66]. To account for nonreciprocal interactions, we instead assume that each spin tends to minimize its own selfish energy , and choose the probability of flipping at each time step to be given by the Glauber rule
@@
where is the change in selfish energy between both configurations, is the temperature, and the Boltzmann constant. Since we define different potential functions for each spin, the dynamics cannot be derived from a single potential function and detailed balance is broken. This formulation encompasses several nonreciprocal dynamics used among the physical sciences [67, 68, 69, 70, 71, 72, 73, 74, 75, 76].
@@
The nonreciprocal Ising model— Systems with nonreciprocal couplings do not always exhibit macroscopic oscillations. These tend to occur, when a small number of species have competing goals. Correspondingly, we split the spins into two species, labeled by Greek indices , located on sites of a -dimensional cubic lattice of linear size (Fig. 1b). Spins of the same species tend to align with their neighbors, while spins of different species interact in a nonreciprocal manner: spins tend to align with spins , whereas spins tend to anti-align with spins . This is captured by the selfish energy
@@
with , where the first term captures the intra-species interactions with the sum running over nearest neighbors of , and the second term captures the inter-species interactions where summation over species indices is implied ( is the Levi-Civita symbol). We have assumed for simplicity that inter-species interactions only occur on-site and are purely nonreciprocal (general asymmetric on-site interactions are considered in [77]).
@@
Mean-field equation— Within the mean-field approximation and upon rescaling of time and space, the dynamics of the average magnetization is [77]
@@
The phase diagram in Fig. 1c shows the stable homogeneous solutions of Eq. (3) as a function of the couplings and . When nonreciprocal interactions are turned off (), a pitchfork bifurcation at (red point in Fig. 1c) separates a disordered phase (in blue) from a ferromagnetic phase (in green), as in the equilibrium Ising model. When nonreciprocal interactions are present (), a time-dependent oscillatory state that we dub “swap phase” arises (in pink). This limit cycle state where both species flip their magnetizations repeatedly is separated from the disordered state by a Hopf bifurcation at (green line), and from the ferromagnetic phase by another bifurcation known as a saddle-node on an invariant circle (SNIC) bifurcation [78, 79] (blue line). The shape of the limit cycle in () space evolves from a circle near the Hopf bifurcation to a square near the SNIC bifurcation [77]. The oscillation period remains finite at the Hopf bifurcation, and it diverges at the SNIC bifurcation [78].
Monte-Carlo simulations— To go beyond the mean-field description, we perform large-scale Monte-Carlo (MC) simulations of the nonreciprocal Ising model in both two and three dimensions (see [77] for details on the update rules; we did not observe qualitative differences with other update rules). Unless mentioned otherwise, we initialize the system in an ordered state where all spins of the same species are either up or down, and let the system evolve until a steady-state is reached.
The qualitative results of our simulations are summarized in Fig. 1d-e. In 2D, any amount of nonreciprocity destroys all order. In 3D, our simulations suggest that the swap phase survives the thermodynamic limit, while the static ferromagnetic phase is eventually destroyed by any amount of nonreciprocity.
@@
and the phase space angular momentum related to entropy production rate [80, 77]
where is the angle on the plane (Fig. 2a), are the total magnetizations and the average is over time and realizations. The synchronization order parameter is zero if the system is disordered and nonzero in both the swap and static-order phases, while is zero in the disordered and static-order phases and non-zero in the swap phase [77].
Destruction of the swap phase in 2D by spiral defects— In small 2D systems, simulations show states in which and oscillate in time (like in the mean-field swap phase) with a large amplitude. However, oscillations are irregular, and typically occur through nucleation of droplets of opposite magnetization by the “unsatisfied” species, in alternating order (see Movie 1). As the size increases, it becomes apparent that this oscillatory state is a transient, eventually destabilized by proliferation of spiral defects similar to the rotational-symmetric spirals observed in the complex Ginzburg-Landau equation [82, 83, 84, 47, 85, 86, 87, 88, 89, 90, 91, 92, 93], but whose shape has 4-fold rotational symmetry (four “arms”). This is evidenced in Fig. 2b-c, which shows the emergence of defects in the angle and the resulting drop in the magnetizations and after an oscillatory transient, along with snapshots of at different stages (see also Movie 1).
@@
Existence of a stable swap phase in 3D and critical exponents— In 3D, we find that a stable swap phase does exist. Its behavior is demonstrated in Fig. 3a where we show a kymograph of the -field of a single 1D row in the 3D lattice as well as plot the total magnetizations as a function of time (see also Movie 2). The oscillations are spatially homogeneous and and have a fixed period and phase shift. The coherence time of and diverges with system size as (Fig. 3b), in agreement with a temporal crystal behavior (Fig. 1a) [94]. In the thermodynamic limit, coherent oscillations persist indefinitely despite the presence of noise.
@@
Figure 3c shows that the phase-space trajectories become circular when approaching the phase transition line, supporting a Hopf behavior [95]. Renormalization group studies based on the -expansion suggest that the universality class associated to the Hopf bifurcation is formally equivalent to the universality class of the ferromagnet/paramagnet transition of the XY model [96, 97]. Using finite-size scaling [77] and setting we determine the critical exponents of the correlation length, susceptibility, and order parameter: , , and , which are indeed in good agreement with 3D XY critical exponents [98, 99], more so than with 3D Ising critical exponents [98] corresponding to the case, see Fig. 3f.
For a fixed , the critical separating disorder from swap depends on , in contrast with the mean-field prediction, in which is independent of (Fig. 1c). Above some critical value, the phase transition loses its second-order-like nature (see region in Fig. 3d where becomes discontinuous), apparently due to scroll waves (3D analog of spiral waves) [100, 101, 102, 103, 104] that destabilize the swap phase [77]. Within the size limitations of our simulations, it is unclear what part of the swap phase is destroyed by scroll waves in the thermodynamic limit and what type of phase transition is associated with their appearance (Fig. 1e).
@@
Instability of the static-order phase due to droplet growth in any finite dimension— In finite systems, a ferromagnetic-like state with static order is observed both in 2D and 3D. As system size increases, however, this static state is destabilized by nucleation of droplets (see Fig. 4a and Movie 4) that grow and flip the magnetization, in alternating order of A- and B-spins. As a result, the static-order phase is replaced by droplet-induced oscillations (different in nature from the homogeneous noisy oscillations close to the Hopf bifurcation shown in Fig. 3a and Movie 2) with oscillation period that converges to a finite value as system size increases (Fig 4b). This is further supported by Figs. 2e and 3e (right parts) showing that the -transition-point between static order and oscillations in finite systems (indicated by a sharp change in the slope of ) increases with size.
To see why nonreciprocity allows droplets to grow, in contrast with the equilibrium Ising model where they tend to shrink, assume first that a system with small nonreciprocal coupling is in the static-order phase, so most spins in both lattices are up [105]. Since the system is static, it can be mapped into two equilibrium Ising models (Fig. 4c) with opposite magnetic fields, with energy where for A-spins and for B-spins (see Eq. (2)). While sub-system A is in a stable state in the effective equilibrium system, sub-system B is in a metastable state, as it prefers a state with opposite magnetization. For large enough system size, sub-system then transitions to its equilibrium stable state by nucleating droplets larger than a critical size whose value depends on and [106, 105]. After a droplet has expanded beyond , the stable sub-system becomes metastable and nucleates droplets, and so on (Fig. 4c). Crucially, for finite and , is finite, making the static-order phase unstable in the thermodynamic limit in any finite dimension. Note that when the symmetry between the spin species is broken, droplets do not expand at the same velocity, leading to droplet-capture and shrinkage that can re-stabilize the static-order phase (Fig. 4d and Movie 5). This can be achieved, for example, by modifying the inter-species interaction in the selfish energy, Eq. (2), to be where and (see Ref. [77] for details).
@@
We have shown that a nonreciprocal generalization of the 3D Ising model can act as a stable spatially-distributed clock characterized by well-defined critical properties reminiscent of time-crystals [113, 114, 115, 100, 116, 117, 118, 119, 120, 121]. While minimalistic, this model contains features arising in models of the human brain [67], opinion dynamics [122, 123], spinor BECs [124, 125, 31] and micromechanical oscillators [69], which can all be modeled by non-potential spin systems.
Code availability — The code used for performing the Monte-Carlo simulations, computing the critical exponents, and plotting the mean-field phase portrait is available under the 2-clause BSD license at https://doi.org/10.5281/zenodo.14816551
@@ Acknowledgements.
- [81] The sign of the phase space angular momentum in Eq. (5) is chosen such that it is positive when is positive.
@@
- [92] Yiwei Zhang and Étienne Fodor. Pulsating active matter. Physical Review Letters, 131(23):238302, 2023.
- [93] Alessandro Manacorda and Étienne Fodor. Pulsating with discrete symmetry. arXiv preprint arXiv:2310.14370, 2023.
@@
- [95] Yoshiki Kuramoto. Chemical Oscillations, Waves, and Turbulence. Springer, 1984.
@@
- [120] Phatthamon Kongkhambut, Jim Skulte, Ludwig Mathey, Jayson G Cosme, Andreas Hemmerich, and Hans Keßler. Observation of a continuous time crystal. Science, 377(6606):670–673, 2022.
- [121] Xiaoling Wu, Zhuqing Wang, Fan Yang, Ruochen Gao, Chao Liang, Meng Khoon Tey, Xiangliang Li, Thomas Pohl, and Li You. Dissipative time crystal in a strongly interacting rydberg gas. Nature Physics, 20(9):1389–1394, 2024.